(x^2+1) y"+xy'-y=0 solve by power series
1
Expert's answer
2020-04-22T18:55:47-0400

Problem:

"(x^2+1)\times \dfrac{\partial^2 y}{\partial x^2}+x\times \dfrac{\partial y}{\partial x}-y=0" (1)

Let's image solution of the task in the next view:

"y=x^\sigma(a_{0}+a_{1}x+a_{2}x^2+...a_{k}x^k+...)" (2)

Now subsitute equation (2) into equation (1):

"(x^2+1)\times(\sigma\times(\sigma-1)\times a_{0}x^{\sigma-2}+(\sigma+1)\times\sigma\times a_{1}x^{\sigma-1}+(\sigma+2)\times(\sigma+1)\times a_{2}x^{\sigma}+...)+x\times(\sigma\times a_{0}x^{\sigma-1}+(\sigma+1)\times a_{1}x^{\sigma}+(\sigma+2)\times a_{2}x^{\sigma+1}+...)-(a_{0}x^{\sigma}+a_{1}x^{\sigma+1}+a_{2}x^{\sigma+2}+...)=0"

Or:

"\sigma\times(\sigma-1)\times a_{0}x^{\sigma-2}
+ (\sigma+1)\times\sigma\times a_{1}x^{\sigma-1} + 
(\sigma\times(\sigma-1)\times a_{0}+(\sigma+2)\times(\sigma+1)\times a_{2}+\sigma\times a_{0}-a_{0})x^{\sigma}+...=0"

Now equate all coefficient near "x^{\sigma-2}" , "x^{\sigma-1}", "x^{\sigma}",... to zero:

"\begin{cases}
a_{0}\times(\sigma-1)\times\sigma=0 \\
a_{1}\times(\sigma+1)\times\sigma=0 \\
a_{0}\times(\sigma+1)\times(\sigma-1)-a_{2}\times(\sigma+2)\times(\sigma+1)=0 \\
a_{1}\times\sigma\times(\sigma+2)-a_{3}\times(\sigma+3)\times(\sigma+2)=0\\
.... \\
a_{k}\times(\sigma-k+1)\times(\sigma+k+1)-a_{k+2}\times(\sigma+k+2)\times(\sigma+k+1)=0\\
\end{cases}" (3)

From the first equation (using that "a_{0}\not=0") we get:

"\sigma=0" or "\sigma=1"

Case: "\sigma=1"

Our system (3) will have the next view:

"\begin{cases}
0=0\\
a_{1}\times 2=0\\
a_{0}\times 0-a_{2}\times 3=0\\
a_{1}\times 3-a_{3}\times12=0\\
...
\end{cases}"

Or:

"\begin{cases}
0=0\\
a_{1}=0\\
a_{2}=0\\
a_{3}=0\\
...\\
\end{cases}"

So we have that all coefficient "a_{k}" (except "a_{0}") are equal zero. So solution of this problem in this case is:

"y=a_{0}\times x^{\sigma}",

But for this case "\sigma=1", so

"y=a_{0}\times x", where "a_{0}" is a different constant.

Case: "\sigma=0"

Our system (3) will have the next view:

"\begin{cases}
0=0\\
0=0\\
-a_{0}-2\times a_{2}=0\\
a_{1}\times 0 - a_{3}\times 6=0\\
...
\end{cases}"

Or:

"\begin{cases}
a_{2}=-\dfrac{a_{0}}{2}\\
a_{3}=0\\
...
\end{cases}"

So we have that coefficient "a_{3}", "a_{5}", ... are equals to zero, and the pairs coefficient(and "a_{1}") aren't equal to zero. For convenience let's get: "a_{1}=0".

For this case, the genral solution will have the next view:

"y=a_{0}+a_{2} x^2+ a_{4}x^4+...=
\begin{matrix}
 \\
\sum \\
_{k=0}
\end{matrix}
a_{2k}x^{2k}"